# RILe: Reinforced Imitation Learning

### v2

@@ Abstract
Reinforcement Learning has achieved significant success in generating complex behavior but often requires extensive reward function engineering. Adversarial variants of Imitation Learning and Inverse Reinforcement Learning offer an alternative by learning policies from expert demonstrations via a discriminator. However, these methods struggle in complex tasks where randomly sampling expert-like behaviors is challenging. This limitation stems from their reliance on policy-agnostic discriminators, which provide insufficient guidance for agent improvement, especially as task complexity increases and expert behavior becomes more distinct. We introduce RILe (Reinforced Imitation Learning environment), a novel trainer-student system that learns a dynamic reward function based on the student’s performance and alignment with expert demonstrations. In RILe, the student learns an action policy while the trainer, using reinforcement learning, continuously updates itself via the discriminator’s feedback to optimize the alignment between the student and the expert. The trainer optimizes for long-term cumulative rewards from the discriminator, enabling it to provide nuanced feedback that accounts for the complexity of the task and the student’s current capabilities. This approach allows for greater exploration of agent actions by providing graduated feedback rather than binary expert/non-expert classifications. By reducing dependence on policy-agnostic discriminators, RILe enables better performance in complex settings where traditional methods falter, outperforming existing methods by 2x in complex simulated robot-locomotion tasks.
@@ 1 Introduction
Reinforcement Learning (RL) has emerged as a powerful framework for teaching agents to perform complex tasks. In recent years, deep reinforcement learning has demonstrated remarkable success in replicating sophisticated behaviors, including playing Atari games, chess, and Go [1, 2]. However, these achievements often come at a cost: the tedious and challenging process of designing reward functions, as predicting the policy outcome from a manually crafted reward function remains notoriously difficult.
To overcome the reward engineering problem, Imitation Learning (IL) leverages expert demonstrations to learn a policy. Since vast amounts of expert data are required to accurately learn expert behaviors, Adversarial Imitation Learning (AIL) approaches, such as GAIL [3], have been proposed as data-efficient alternatives. AIL employs a discriminator to measure similarity between learned behavior and expert behavior, rewarding the agent accordingly. While computationally efficient, AIL methods suffer from a critical limitation: the policy-agnostic nature of their discriminators. The discriminator lacks any inherent incentive to guide the agent towards expert-like behavior, in contrast to engineered reward functions in RL. Consequently, AIL methods face challenges in complex tasks requiring extensive exploration to find optimal actions. For instance, in digital locomotion tasks, AIL methods often struggle to consistently replicate expert performance [4].
Inverse Reinforcement Learning (IRL) is another approach to alleviate reward engineering. Unlike IL, which directly learns expert behavior, IRL seeks to infer the underlying reward function that motivates the agent to acquire expert behaviors. The reward function and the agent are trained iteratively, with updates to the reward function based on the agent’s behavior. This iterative process renders IRL computationally expensive [5]. Adversarial Inverse Reinforcement Learning (AIRL) [6] attempts to address this inefficiency by introducing a discriminator that enables simultaneous learning of the policy and reward function. However, in AIRL, the reward function is tightly coupled to the discriminator, potentially limiting its ability to capture complex task structures or long-term dependencies and inheriting the limitations of a policy-agnostic discriminators. This highlights the need for a method that can learn a more flexible reward function without the computational overhead of traditional IRL methods.
To overcome these challenges and effectively learn behaviors in complex settings, we propose Reinforced Imitation Learning (RILe) (Fig. 1-(d)). RILe aims to combine the ability to learn a reward function that actively guides the agent to imitate expert behavior with the computational efficiency of adversarial frameworks. At the core of RILe is a novel trainer-student system designed to address the shortcomings of existing methods:
  • A student agent that learns to replicate the expert’s policy via RL in the environment
  • A trainer agent that learns a reward function via RL and guides the student agent during training
By integrating the trainer-student dynamic, RILe decouples reward learning from policy learning and the discriminator, allowing each component to specialize and thereby overcome the limitations of policy-agnostic discriminators. While RILe utilizes a discriminator similar to those in adversarial frameworks, its role is fundamentally redefined. In RILe, the discriminator’s primary function is to provide feedback to the trainer agent by distinguishing expert data from student roll-outs. This feedback serves as the reward signal for the trainer, not directly influencing the student agent. The trainer leverages the discriminator’s feedback to learn a reward function that effectively guides the student agent. This approach enables more nuanced reward shaping, particularly beneficial in tasks requiring complex decision-making and extensive exploration.
@@
    Decoupled Reward-function Learning: We introduce a novel approach where the trainer agent learns the reward function independently from both the student agent and the discriminator. Unlike existing methods that derive rewards directly from discriminator outputs, our trainer agent uses reinforcement learning to optimize the reward function based on the feedback from the discriminator. By focusing on long-term reward maximization, RL enables the trainer to distill inconsistent feedback from the discriminator into meaningful rewards, leading to better student performance.
@@
    Dynamic Reward Customization: Our trainer agent dynamically adjusts rewards based on the student agent’s progress, facilitating a better learning experience and enabling accurate imitation of expert behavior in complex settings. This adaptive approach allows for more gradual learning, particularly in tasks where the optimal behavior may change depending on the agent’s current capabilities.
We evaluate RILe against state-of-the-art methods in AIL, and AIRL, specifically GAIL [3], and AIRL [6]. Our experiments span three distinct scenarios: (1) Tailoring a reward function dynamically in a discrete maze task, (2) Investigating the impact of expert data on the trainer-student dynamics in a humanoid locomotion task, and (3) Imitating motion-capture data in continuous control tasks. The results demonstrate RILe’s superior performance, especially in complex tasks, and its ability to learn an effective dynamic reward function where baseline methods fail.
@@ 2 Related Work
We review literature on learning from expert demonstrations, focusing on Imitation Learning (IL) and Inverse Reinforcement Learning (IRL), which form the conceptual foundation of RILe.
@@
Early work introduced Behavioral Cloning (BC) [7], which learns a policy congruent with expert demonstrations through supervised learning. DAgger [8] introduces data aggregation. GAIL [3] introduces adversarial methods, where a discriminator aims to discriminate expert demonstrations, while a generator tries to fool the discriminator. DQfD [9] proposes two-stage approach with pre-training, and ValueDice [10] uses a distribution-matching objective between policy and expert. Despite progress, IL faces challenges in efficacy and generalization [5, 11]. RILe addresses these by introducing an adaptive teacher agent to guide the student beyond expert demonstrations.
@@
IRL, introduced by [12], learns the expert’s intrinsic reward function. Key developments include Apprenticeship Learning [13], Maximum Entropy IRL [14], and adversarial approaches like AIRL [6]. Recent work explores handling unstructured data [15] and cross-embodiment scenarios [16]. Despite advancements, IRL faces challenges in computational efficiency and scalability [17]. RILe addresses these by jointly learning policy and reward function in a single process.
@@ 3.1 Markov Decision Process
A standard Markov Decision Process (MDP) is defined by $(S,A,R,T,K,\gamma)$. $S$ is the state space that consists of all possible environment states $s$, and $A$ is action space that contains all possible environment actions $a$. $R=R(s,a):S\ \times\ A\ \rightarrow\ \mathbb{R}$ is the reward function. $T=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is defined as an unknown state state transition probability function upon taking action $a\in A$ in state $s\in S$. $K(s)$ is the initial state distribution, i.e., $s_{0}\sim K(s)$ and $\gamma$ is the discount factor. The policy $\pi=\pi(a|s):S\ \rightarrow\ A$ is a mapping from states to actions. In this work, we consider $\gamma$-discounted infinite horizon settings. Following [3], expectation with respect to the policy $\pi\in\Pi$ refers to the expectation when actions are sampled from $\pi(s)$: $E_{\pi}[R(s,a)]\triangleq E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$, where $s_{0}$ is sampled from an initial state distribution $\text{K}(s)$, $a_{t}$ is given by $\pi(\cdot|s_{t})$ and $s_{t+1}$ is determined by the unknown transition model as $P(\cdot|s_{t},a_{t})$. The unknown reward function $R(s,a)$ generates a reward given a state-action pair $(s,a)$. We consider a setting where $R=R(s,a)$ is parameterized by $\theta$ as $R_{\theta}(s,a)\in\mathbb{R}$ [18].
Our work considers an imitation learning problem from expert trajectories, each consisting of states $s$ and actions $a$. The set of expert trajectories $\tau_{E}$ are sampled from an expert policy $\pi_{E}\in\Pi$, where $\Pi$ is the set of all possible policies. We assume that we have access to $m$ expert trajectories, all of which have $n$ time-steps, $\tau_{E}=\{(s^{i}_{0},a^{i}_{0}),(s^{i}_{1},a^{i}_{1}),\ldots,(s^{i}_{n},a^{i}_{n})\}_{i=1}^{m}$.
@@ 3.2 Reinforcement Learning (RL)
Reinforcement learning seeks to find an optimal policy, $\pi^{*}$. that maximizes the discounted cumulative reward given from the reward function $R=R(s,a)$ (Fig. 1-(a)). In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=E_{\pi}[-\text{log}\ \pi(a|s)]$ [3, 19]. The RL problem with a parameterized reward function and entropy regularization is defined as
$\text{RL}(R_{\theta}(s,a))=\pi^{*}=\operatorname*{arg\,max}_{\pi}E_{\pi}[R_{\theta}(s,a)]+H(\pi).$ (1)
@@
Figure 1: Overview of the related works. (a) Reinforcement Learning (RL): learning a policy that maximizes hand-defined reward function; (b) Inverse RL (IRL): learning a reward function from data. IRL has two stages: 1. training a policy with frozen reward function, and 2. updating the reward function by comparing the converged policy with data. These stages repeated several times; (C) Generative Adversarial Imitation Learning (GAIL) + Adversarial IRL (AIRL): using discriminator as a reward function. GAIL trains both policy and the discriminator at the same time. AIRL implements a new structure on the discriminator, seperating reward from environment dynamics by using two networks under the discriminator (see additional terms in green). (D) RILe: similar to IRL, learning a reward function from data. RILe learns the reward function at the same time with the policy, using discriminator as a guide for learning the reward.
@@ 3.3 Inverse Reinforcement Learning (IRL)
Given sample trajectories $\tau_{E}$ from an optimal expert policy $\pi_{E}$, inverse reinforcement learning aims to recover a reward function $R^{*}_{\theta}(s,a)$ that maximally rewards the expert’s behavior (Fig. 1-(b)). Formally, IRL seeks a reward function, $R_{\theta}^{*}(s,a)$, satisfying: $E_{\pi_{E}}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\geq E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\quad\forall\pi.$ Optimizing this reward function with reinforcement learning yields a policy that replicates expert behavior: $\text{RL}(R_{\theta}^{*}(s,a))=\pi^{*}=\operatorname*{arg\,max}_{\pi}E_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]$.
Since only the expert’s trajectories are observed, expectations over $\pi_{E}$ are estimated from samples in $\tau_{E}$. Incorporating entropy regularization $H(\pi)$, maximum causal entropy inverse reinforcement learning [14] is defined as
$\text{IRL}(\tau_{E})=\operatorname*{arg\,max}_{R_{\theta}(s,a)\in\mathbb{R}}\left(E_{s,a\in\tau_{E}}[R_{\theta}(s,a)]-\max_{\pi}\left(E_{\pi}[R_{\theta}(s,a)]+H(\pi)\right)\right).$ (2)

### v3

@@ Abstract
Acquiring complex behaviors is essential for artificially intelligent agents, yet learning these behaviors in high-dimensional settings poses a significant challenge due to the vast search space. Traditional reinforcement learning (RL) requires extensive manual effort for reward function engineering. Inverse reinforcement learning (IRL) uncovers reward functions from expert demonstrations but relies on an iterative process that is often computationally expensive. Imitation learning (IL) provides a more efficient alternative by directly comparing an agent’s actions to expert demonstrations; however, in high-dimensional environments, such direct comparisons offer insufficient feedback for effective learning. We introduce RILe (Reinforced Imitation Learning), a framework that combines the strengths of imitation learning and inverse reinforcement learning to learn a dense reward function efficiently and achieve strong performance in high-dimensional tasks. RILe employs a novel trainer–student framework: the trainer learns an adaptive reward function, and the student uses this reward signal to imitate expert behaviors. By dynamically adjusting its guidance as the student evolves, the trainer provides nuanced feedback across different phases of learning. Our framework produces high-performing policies in high-dimensional tasks where direct imitation fails to replicate complex behaviors. We validate RILe in challenging robotic locomotion tasks, demonstrating that it significantly outperforms existing methods and achieves near-expert performance across multiple settings.
@@ 1 Introduction
Learning complex behaviors is critical for advancing artificially intelligent agents in fields such as robotics and strategic games. Over the years, reinforcement learning (RL) has emerged as a powerful framework for teaching agents to perform sophisticated tasks, yet it often requires extensive manual reward function design, which is both time-consuming and error-prone.
There are two ways to address the reward engineering problem. First, Inverse Reinforcement Learning (IRL) [1, 2] offers a remedy by inferring the reward function from expert demonstrations, thus reducing the burden of manual reward engineering. IRL proceeds iteratively: it first trains a policy (the learning agent’s decision-making mechanism) using the current reward function, observes how well the agent’s behavior aligns with the expert’s, and then refines the reward function to better guide the policy toward expert-like behaviors. Repeating this process eventually yields a reward function capable of providing nuanced feedback at different stages of learning. However, this iterative procedure is computationally expensive [3], especially in high-dimensional environments where both the reward and the policy must explore a large state-action space.
Second, Imitation learning (IL) bypasses explicit reward design by directly comparing learned behaviors to expert demonstrations via a comparison mechanism. Traditional IL approaches such as Behavioral Cloning (BC) [4] match the learned actions to expert demonstrations directly, requiring a substantial amount of expert data in high-dimensional tasks. To improve data efficiency, Adversarial Imitation Learning (AIL) methods, such as GAIL [5], introduce a discriminator that operates as a comparison mechanism and judges how expert-like the learned behaviors are. However, both traditional IL and AIL lack a reward function that emphasizes specific subgoals or partial improvements. Instead, they rely on a comparison metric, often a distance measure or binary classification, that merely checks whether the agent’s behavior is (or is not) similar to the expert. Such comparison-based signals offer no fine-grained guidance on which specific actions or sub-strategies to prioritize. Consequently, both traditional IL and AIL struggle in high-dimensional environments [6, 7], where the agent needs more granular and adaptive feedback than these mechanisms provide.
Adversarial Inverse Reinforcement Learning (AIRL) [8] attempts to remedy IRL’s inefficiency by integrating a learned reward function within a discriminator. However, AIRL tightly couples the reward function to the discriminator’s output, causing it to inherit AIL’s limitations in high-dimensional settings where more fine-grained guidance is needed.
Real-life learning scenarios suggest a different approach: think of parents and children, or a pet owner and their dog. The teacher also refines how they teach as the student progresses. Each success or failure in the student’s understanding informs the teacher’s approach, creating a positive-sum relationship: lessons learned from suboptimal behaviors ultimately yield better trainers, which, in turn, guide the student’s learning progress more effectively. By contrast, existing approaches lack this continuous cooperative synergy. Adversarial Imitation Learning (AIL) does update a discriminator alongside the policy, but the discriminator’s sole role is to distinguish expert-like behavior from non-expert behavior. Consequently, the student attempts to fool this judge into classifying its behavior as expert-like, resulting in a competitive process rather than a cooperative trainer that dynamically shapes rewards based on suboptimal behaviors. Meanwhile, IRL methods only refine the reward after the policy converges, missing the opportunity for continuous co-evolution throughout training.
To address these issues, inspiring by these insights, we propose Reinforced Imitation Learning (RILe). RIle combines the adaptive reward benefits of IRL with the computational efficiency of AIL (Fig. 1-(d)). RILe is a novel trainer-student system that establishes a positive-sum relationship between the trainer and the student. Specifically, RILe is composed of:
  • Student Agent: Learns a policy to imitate expert demonstrations using reinforcement learning.
  • Trainer Agent: Simultaneously learns a reward function using reinforcement learning, leveraging an adversarial discriminator for continuous feedback on student performance.
RILe’s trainer continuously updates the reward function in tandem with the student’s policy updates, where IRL refines its reward function only after training a policy to convergence. Specifically, the trainer queries a discriminator to measure how expert-like the student’s behavior is, then optimizes the reward function based on that feedback, without waiting for the policy to converge. RILe preserves nuanced reward shaping, while avoiding IRL’s heavy computational loop. As a result, RILe is particularly effective in high-dimensional settings, where agents need fine-grained guidance at every stage of learning.
@@
    Efficient Reward-Function Learning via RL: We introduce a reinforcement-learning-based approach for training a reward function simultaneously with the policy. This avoids IRL’s repeated policy re-training and the purely discriminator-based rewards of AIL/AIRL. Reinforcement learning enables the trainer agent to explore multiple reward regimes, yielding a reward function that considers long-horizon effects.
@@
    Dynamic Reward Customization: RILe offers context-sensitive guidance at every stage of training, because the trainer agent updates the reward function as the student evolves. This dynamic shaping is especially valuable in high-dimensional tasks, where the learning agent requires different forms of encouragement during intermediate-stages than later-stages of the training. Consequently, RILe enables accurate imitation of the expert performance in high-dimensional tasks.
We evaluate RILe in comparison to state-of-the-art methods in AIL, IRL, and AIRL, specifically GAIL [5] AIRL [8], GAIfO [9], BCO [10], IQ-Learn [7] and DRAIL [11]. Our experiments spans four scenarios: (1) Empirically investigating how RILe’s reward-learning strategy differs from AIL and AIRL in a maze task, (2) Quantifying the adaptability of the learned reward function in a humanoid locomotion task, (3) Assessing RILe’s performance in both low- and high-dimensional continuous-control problems, and (4) Analyzing the impact of explicit expert-data usage within RILe’s training process. Our results show RILe’s superior performance, particularly in high-dimensional environments, and highlight RILe’s ability to learn a dynamic reward function that effectively guides the student through multiple stages of training.
@@ 2 Related Work
We review research on learning from expert demonstrations, focusing on Imitation Learning (IL) and Inverse Reinforcement Learning (IRL), the conceptual foundations of RILe.
@@
Early work in IL introduced Behavioral Cloning (BC) [4], which frames policy learning as a supervised problem where the agent’s actions are directly matched to expert demonstrations. DAgger [12] refines BC by aggregating data over multiple iterations to mitigate compounding errors. GAIL [5] employs adversarial training: a discriminator learns to distinguish expert trajectories from the agent’s, while the generator (agent) adapts to mimic expert-like behavior. BCO [10] extends BC, and GAIfO [9] extends GAIL, both to handle state-only observation scenarios. DQfD [13] introduces a two-stage approach with pre-training, while ValueDice [14] aligns policy and expert distributions via a distribution-matching objective. More recently, DRAIL [11] leverages a diffusion-based discriminator to enhance learning efficiency in adversarial imitation.
Despite these advances, IL methods face challenges in high-dimensional environments [6, 7], where naive action matching or purely adversarial comparisons fail to provide sufficiently granular guidance. RILe addresses these limitations through an adaptive trainer–student framework, where a learned reward function provides more nuanced guidance than standard IL comparison mechanisms.
@@
Inverse Reinforcement Learning (IRL), introduced by [1], aims to uncover the expert’s intrinsic reward function from demonstrations. Major developments include Apprenticeship Learning [15], Maximum Entropy IRL [2], and adversarial variants like AIRL [8]. IQ-Learn [7] reformulates IRL by integrating the inverse reward learning process into Q-learning for better scalability. More recent work focuses on unstructured data [16] and cross-embodiment transfer [17].
Nonetheless, IRL methods struggle with computational inefficiency and limited scalability [18], particularly in high-dimensional tasks where repeated iterations of policy learning and reward refinement become costly. RILe mitigates these challenges by jointly learning the policy and reward function in a single process, avoiding IRL’s iterative retraining loop and facilitating more efficient reward shaping for complex environments.
@@ 3.1 Markov Decision Process
A standard Markov Decision Process (MDP) is defined by $(S,A,R,T,K,\gamma)$. $S$ is the state space consisting of all possible environment states $s$, and $A$ is action space containing all possible environment actions $a$. $R=R(s,a):S\ \times\ A\ \rightarrow\ \mathbb{R}$ is the reward function. $T=\{P(\cdot|s,a)\}$ is the transition dynamics where $P(\cdot|s,a)$ is an unknown state state transition probability function upon taking action $a\in A$ in state $s\in S$. $K(s)$ is the initial state distribution, i.e., $s_{0}\sim K(s)$ and $\gamma$ is the discount factor. The policy $\pi=\pi(a|s):S\ \rightarrow\ A$ is a mapping from states to actions. In this work, we consider $\gamma$-discounted infinite horizon settings. Following [5], expectation with respect to the policy $\pi\in\Pi$ refers to the expectation when actions are sampled from $\pi(s)$: $\mathbb{E}_{\pi}[R(s,a)]\triangleq\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t})]$, where $s_{0}$ is sampled from an initial state distribution $\text{K}(s)$, $a_{t}$ is given by $\pi(\cdot|s_{t})$ and $s_{t+1}$ is determined by the unknown transition model as $P(\cdot|s_{t},a_{t})$. The unknown reward function $R(s,a)$ generates a reward given a state-action pair $(s,a)$. We consider a setting where $R=R(s,a)$ is parameterized by $\theta$ as $R_{\theta}(s,a)\in\mathbb{R}$ [19].
Our work considers an imitation learning problem from expert trajectories, consisting of states $s$ and actions $a$. The set of expert trajectories $\tau_{E}$ are sampled from an expert policy $\pi_{E}\in\Pi$, where $\Pi$ is the set of all possible policies. We assume that we have access to $m$ expert trajectories, all of which have $n$ time-steps, $\tau_{E}=\{(s^{i}_{0},a^{i}_{0}),(s^{i}_{1},a^{i}_{1}),\ldots,(s^{i}_{n},a^{i}_{n})\}_{i=1}^{m}$.
@@ 3.2 Reinforcement Learning (RL)
Reinforcement learning seeks to find an optimal policy, $\pi^{*}$. that maximizes the discounted cumulative reward given from the reward function $R=R(s,a)$ (Fig. 1-(a)). In this work, we incorporate entropy regularization using the $\gamma$-discounted casual entropy function $H(\pi)=\mathbb{E}_{\pi}[-\text{log}\ \pi(a|s)]$ [5, 20]. The RL problem with a parameterized reward function and entropy regularization is defined as:
$\text{RL}(R_{\theta}(s,a))=\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}[R_{\theta}(s,a)]+H(\pi).$ (1)
@@
Figure 1: Overview of the related works. (a) Reinforcement Learning (RL): learning a policy that maximizes hand-defined reward function; (b) Inverse RL (IRL): learning a reward function from data. IRL has two stages: 1. training a policy with frozen reward function, and 2. updating the reward function by comparing the converged policy with data. These stages repeated several times; (C) Generative Adversarial Imitation Learning (GAIL) + Adversarial IRL (AIRL): using discriminator as a reward function. GAIL trains both policy and the discriminator at the same time. AIRL implements a new structure on the discriminator, seperating reward from environment dynamics by using two networks under the discriminator (see additional terms in green). (D) RILe: similar to IRL, learning a reward function from data. RILe learns the reward function at the same time with the policy, using a discriminator as a guide for learning the reward.
@@ 3.3 Inverse Reinforcement Learning (IRL)
Given sample trajectories $\tau_{E}$ from an optimal expert policy $\pi_{E}$, inverse reinforcement learning aims to recover a reward function $R^{*}_{\theta}(s,a)$ that maximally rewards the expert’s behavior (Fig. 1-(b)). Formally, IRL seeks a reward function, $R_{\theta}^{*}(s,a)$, satisfying: $\mathbb{E}_{\pi_{E}}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})]\geq\mathbb{E}_{\pi}[\sum_{t=0}^{\infty}\gamma^{t}R_{\theta}^{*}(s_{t},a_{t})+H(\pi)]\quad\forall\pi.$ Optimizing this reward function with reinforcement learning yields a policy that replicates expert behavior: $\text{RL}(R_{\theta}^{*}(s,a))=\pi^{*}$. Since only the expert’s trajectories are observed, expectations over $\pi_{E}$ are estimated from samples in $\tau_{E}$. Incorporating entropy regularization $H(\pi)$, maximum causal entropy inverse reinforcement learning [2] is defined as
$\text{IRL}(\tau_{E})=\operatorname*{arg\,max}_{R_{\theta}(s,a)\in\mathbb{R}}\left(\mathbb{E}_{s,a\in\tau_{E}}[R_{\theta}(s,a)]-\max_{\pi}\left(\mathbb{E}_{\pi}[R_{\theta}(s,a)]+H(\pi)\right)\right).$ (2)